# A note for double Hölder regularity of the hydrodynamic pressure for weak solutions of Euler equations

### v2

@@ 3.2. Estimate for $I_{1}$
In fact, we shall only utilise $M\in C^{0}\left(\overline{\Omega_{\delta/2}\setminus\Omega_{\delta}}\right)$ in the sequel.
@@
Here $C_{11}$ and $C_{12}$ depend on $d$ and $\|M\|_{C^{0}\left(\overline{\Omega_{\delta/2}\setminus\Omega_{\delta}}\right)}$ only. The index $d-1-\frac{d-2}{d-2+2\gamma}\cdot(d-1+2\gamma)$ on the right-most term of (3.2) is strictly positive. The assertion given in Lemma 3.1 follows immediately from the inequalities (3.8) and (3.2). ∎
@@ 2. Denote
Utilising Taylor expansion, triangle inequality, and the estimate (3.3) for the Neumann Green function, we deduce that
@@
We proceed in two steps.
1. The case $|x_{1}-x_{2}|\leq c_{\Omega}\delta^{\beta}$. First, from the equation (3.1) and Lemmas 3.1 &\&& 3.2, we deduce
@@
provided that the condition (3.1) is verified. That is, the above estimate holds when $[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$ and $c_{\Omega}\delta\leq|x_{1}-x_{2}|\leq c_{\Omega}\delta^{\beta}$, where $10<\beta<$ 1 is a fixed number given in (3.7), and $c_{\Omega}>0$ is geometrical. Here, the constant $C$ depends only on $d$, $\gamma$, $\|u\|_{C^{0,\gamma}(\Omega)}$, and the $C^{2}$-geometry of $\Omega$.
@@
Now, in view of the definition of Weierstrass-$\wp$ (see (2.7)), we have =pWeierstrass-$\wp=p$ in $\Omega\setminus{\bf supp}({\varphi^{\delta}})$, where $p$ is the hydrodynamical pressure we need to estimate. But ${\bf supp}({\varphi^{\delta}})\subset\Omega\setminus\Omega_{2\delta}$ while $[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$, so $x_{1},x_{2}\notin{\bf supp}({\varphi^{\delta}})$. We thus obtain the desired estimate (1.2) given in Theorem 1.1, for those $x_{1}$, $x_{2}$ satisfying the condition (3.1).
The restriction $|x_{1}-x_{2}|\geq c_{\Omega}\delta$ in the above arguments can be easily removed. That is, we may easily extend to the case where $|x_{1}-x_{2}|$ is as small as we want. Indeed, for any $x_{1},x_{2}$ such that $[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$ but $|x_{1}-x_{2}|<c_{\Omega}\delta$, let us pick a point $x_{\star}$ such that
$[x_{1},x_{\star}]\cup[x_{\star},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}},$
$c_{\Omega}\delta\leq|x_{1}-x_{\star}|=|x_{\star}-x_{2}|\leq c_{\Omega}\delta^{\beta}.$
For instance, we may choose $x_{\star}$ to be any point on the $(d-2)$-dimensional submanifold
$\partial{\bf B}_{c_{\Omega}\cdot\delta^{({1+\beta})/{2}}}(x_{1})\cap\partial{\bf B}_{c_{\Omega}\cdot\delta^{({1+\beta})/{2}}}(x_{2})\cap\Omega_{10\cdot\delta^{\beta}}.$
The above arguments then yield that
$\left|p(x_{1})-p(x_{\star})\right|\leq C_{0}|x_{1}-x_{\star}|^{2\gamma}\quad\text{and}\quad\left|p(x_{\star})-p(x_{2})\right|\leq C_{0}|x_{2}-x_{\star}|^{2\gamma},$
for a constant $C_{0}$ depending only on $d$, $\gamma$, $\|u\|_{C^{0,\gamma}(\Omega)}$, and the $C^{2}$-geometry of $\Omega$. Thus, from the triangle inequality we deduce that
$\left|p(x_{1})-p(x_{2})\right|\leq 2^{1-2\gamma}C_{0}|x_{1}-x_{2}|^{2\gamma}.$
This proves the estimate (1.2) in Theorem 1.1 for any $x_{1},x_{2}$ such that $[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$ and $|x_{1}-x_{2}|\leq c_{\Omega}\delta^{\beta}$.
2. The general case. Let $x_{1},x_{2}\in\Omega$ be arbitrary. Recall that $\beta=\beta(d,\gamma):=\frac{d-2}{d-2+2\gamma}$ given in (3.7). Let us take
[…]
See condition (2.2) for the dependency on ${\rm injrad}(\Omega)$ of $\delta$. Such a choice of $\delta$ ensures that there is a piecewise affine curve inside $\Omega_{10\cdot\delta^{\beta}}$ whose endpoints are $\{x_{1},x_{2}\}$, and whose nodes $\{z_{0}=x_{1},z_{1},z_{2},\ldots,z_{N}=x_{2}\}$ satisfies
$c_{\Omega}\delta\leq|z_{j}-z_{j-1}|\leq c_{\Omega}\delta^{\beta}\qquad\text{for each }j\in\{1,2,\ldots,N\}$
for the geometrical constant $c_{\Omega}$ as before, as well as
$\sum_{j=1}^{N}|z_{j}-z_{j-1}|\approx|x_{1}-x_{2}|$
modulo another geometrical constant. Indeed, we may first find a geodesic ${\bf c}$ in $\Omega$ connecting $x_{1}$ and $x_{2}$ whose length is comparable to $|x_{1}-x_{2}|$ modulo a geometrical constant, and then construct a polygonal approximation of ${\bf c}$.
With the above preparations, we may apply the arguments in the earlier parts of this subsection to obtain that
$\big{|}p(z_{j})-p(z_{j-1})\big{|}\leq C_{0}|z_{j}-z_{j-1}|^{2\gamma}\qquad\text{for each }j\in\{1,2,\ldots,N\}.$
Again, $C_{0}$ depends only on $d$, $\gamma$, $\|u\|_{C^{0,\gamma}(\Omega)}$, and the $C^{2}$-geometry of $\Omega$. By the triangle and Jensen’s inequalities we have that
$|p(x_{1})-p(x_{2})|$ $\leq\sum_{j=1}^{N}\big{|}p(z_{j})-p(z_{j-1})\big{|}$
$\leq C_{0}\sum_{j=1}^{N}\left(|z_{j}-z_{j-1}|^{2\gamma}\right)$
$\leq C_{0}N^{1-2\gamma}\left(\sum_{j=1}^{N}|z_{j}-z_{j-1}|\right)^{2\gamma}$
$\leq C_{1}N^{1-2\gamma}|x_{1}-x_{2}|^{2\gamma},$ (3.16)
where $C_{1}$ has the same dependency as $C_{0}$ above.
Finally, observe that
$N\leq\frac{\text{ the intrinsic diameter of $\Omega$}\,}{c_{\Omega}\delta^{\beta}}.$ (3.17)
Theorem 1.1 (and the ensuing Remark 1.1) now follows immediately from (3.15), (3.4), and (3.17).
@@ 4. Concluding remarks
In this notes, we have given an alternative, elementary proof for the double Hölder regularity of the Euler pressure $p$ in bounded $C^{2}$-domains in dimension greater than or equal to $3$. We conclude by the following remarks.
@@ 4.1. Two-dimensional case
The proof of Theorem 1.1 above requires $d\geq 3$. In dimension 2, the Neumann Green function satisfies $|D^{\ell}{\mathcal{K}_{N}}(z)|\lesssim|D^{\ell}\log\,z|$ for $\ell\in\{0,1,2\}$, as opposed to the bound (3.3). An adaptation of the arguments in this note should also lead to the same statement of Theorem 1.1 with $d=2$.
4.2. Regularity up to the boundary?
Theorem 1.1 establishes the double Hölder regularity of $p$ in the interior, namely that $p\in C^{0,2\gamma}(\Omega)$, whenever $u\in C^{0,\gamma}(\Omega)$ for bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d\geq 3}$. It has also been established in De Rosa, Latocca and Stefani [13, Theorem 1.1] that if $\Omega$ is a $C^{2,\alpha}$-domain for any $\alpha>0$ and if $u\in C^{0,\gamma}({\Omega};\mathbb{R}^{d})$, then $p$ is continuous up to the boundary; i.e., $p\in C^{0}(\overline{\Omega})$.
However, the approach taken in this note does not yield the double Hölder regularity up to the boundary; that is, we cannot obtain $p\in C^{0,2\gamma}(\overline{\Omega})$. Instead, we may only deduce that $p$ has no “small-scale double Hölder creations” near the boundary in the following sense:
$\limsup_{\kappa\searrow 0}\left(\sup\left\{\frac{|p(x_{1})-p(x_{2})|}{|x_{1}-x_{2}|^{2\gamma}}:\,|x_{1}-x_{2}|\lesssim\min_{i\in\{1,2\}}{\rm dist}(x_{i},{\partial\Omega})\approx\kappa\right\}\right)<\infty,$
where the constants involved in $\lesssim$ and $\approx$ are purely geometrical.
At the moment we are unable to obtain the analogous estimate uniformly in $\kappa$ for $|x_{1}-x_{2}|$ large but $\kappa\searrow 0$. But in this case an upper bound $\mathcal{O}\left(\kappa^{-1+2\gamma}\right)$ for the blowup rate of the double Hölder norm of $p$ has been obtained.
@@ 4.3. Unbounded domains
Acknowledgement. This research was partially supported by NSFC under Grant No. 12331008, and the Shanghai Frontier Research Institute for Modern Analysis. The research of SL was also partially supported by NSFC under Grant Nos. 12201399 and 12411530065, Young Elite Scientists Sponsorship Program by CAST 2023QNRC001, and National Key Research and Development Program 2023YFA1010900. The research of YGW was also partially supported by NSFC under Grant Nos. 12171317, 12250710674 and 12161141004, and Shanghai Municipal Education Commission under Grant No. 2021-01-07-00-02-E00087.
Both authors thank Prof. Claude Bardos, Mr. Daniel Boutros, and Prof. Edriss Titi for pointing out a flawed argument in an earlier version of the manuscript. SL also thanks Professors Linhan Li and Lihe Wang for insightful discussions on Green functions.

### v3

@@ 3.2. Estimate for $I_{1}$
In view of (2.7) and the above argument, we also have ${\mathcal{Q}},{\mathcal{R}},u\otimes u\in C^{0,\gamma}(\Omega)$. Solving the standard Neumann problem from (2.8) and using the definition of ${\mathcal{Q}}$ and ${\mathcal{R}}$, we have $\wp\in C^{0,\gamma}(\Omega)$ and hence
$\|p\|_{C^{0,\gamma}(\Omega)}\leq C\left(d,\gamma,\Omega\right)\|u\|_{C^{0,\gamma}(\Omega)}^{2}.$ (3.7)
In fact, we shall only use $M\in C^{0}\left(\overline{\Omega_{\delta/2}\setminus\Omega_{\delta}}\right)$ and $p\in C^{0}(\Omega)$ in the sequel.
@@
Here $C_{11}$ and $C_{12}$ depend on $d$ and $\|M\|_{C^{0}\left(\overline{\Omega_{\delta/2}\setminus\Omega_{\delta}}\right)}$ only. The index $d-1-\frac{d-2}{d-2+2\gamma}\cdot(d-1+2\gamma)$ on the right-most term of (3.2) is strictly positive. The assertion in Lemma 3.1 now follows from the inequalities (3.9) and (3.2). ∎
@@ 2. Denote
Utilising the Taylor expansion, triangle inequality, and the estimate (3.3) for the Neumann Green function, we deduce that
@@
From (3.1) and Lemmas 3.1 &\&& 3.2, we deduce
@@
provided that the condition (3.1) is verified. That is, the above estimate holds when $[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$ and $c_{\Omega}\delta\leq|x_{1}-x_{2}|\leq c_{\Omega}\delta^{\beta}$, where $10<\beta<$ 1 is a fixed number given in (3.8), and $c_{\Omega}>0$ is geometrical. The constant may be chosen as $C=C(d,\gamma)\cdot C_{\rm geom}\cdot\|u\|_{C^{0,\gamma}(\Omega)}^{2}.$ Then, in view of the definition of Weierstrass-$\wp$ (see (2.7)), we have =pWeierstrass-$\wp=p$ in $\Omega\setminus{\bf supp}({\varphi^{\delta}})$, where $p$ is the pressure we need to estimate. But ${\bf supp}({\varphi^{\delta}})\subset\Omega\setminus\Omega_{2\delta}$ while $[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$, so $x_{1},x_{2}\notin{\bf supp}({\varphi^{\delta}})$. We thus obtain the desired estimate (1.2) in Theorem 1.1, for those $x_{1}$, $x_{2}$ satisfying the condition (3.1).
Now, for any given $x_{1},x_{2}\in\Omega$, we set
$\kappa:=\min\Big{\{}{\rm dist}(x_{1},\partial\Omega),{\rm dist}(x_{2},\partial\Omega)\Big{\}}>0.$
Let $c_{\Omega}$ be the geometrical constant as in the condition (3.1), and consider the two separate cases below:
Case 1: $\rho=|x_{1}-x_{2}|\geq c_{\Omega}(12^{-1}\kappa)^{\frac{1}{\beta}}$ or $\rho\geq\left(c_{\Omega}\right)^{-\frac{\beta}{1-\beta}}$. If the former holds, we deduce from the estimate (3.7) (in fact, we only need the $C^{0}$-bound for $p$; see the comments ensuing this estimate) that
$\frac{\left|p(x_{1})-p(x_{2})\right|}{|x_{1}-x_{2}|^{2\gamma}}$ $\leq 2\cdot 12^{\frac{2\gamma}{\beta}}\|p\|_{C^{0}(\Omega)}\left(c_{\Omega}\right)^{-2\gamma}\kappa^{-\frac{2\gamma}{\beta}}$
$\leq C(d,\gamma,\Omega)\|u\|^{2}_{C^{0,\gamma}(\Omega)}\kappa^{-\frac{2\gamma}{\beta}}.$ (3.16)
If the latter holds, then the analogous argument yields that
$\frac{\left|p(x_{1})-p(x_{2})\right|}{|x_{1}-x_{2}|^{2\gamma}}$ $\leq C(d,\gamma,\Omega)\|u\|^{2}_{C^{0,\gamma}(\Omega)}.$ (3.17)
Case 2: $\rho=|x_{1}-x_{2}|<\min\left\{c_{\Omega}(12^{-1}\kappa)^{\frac{1}{\beta}},\,\left(c_{\Omega}\right)^{-\frac{\beta}{1-\beta}}\right\}$. In this case, taking
$\delta:=\frac{\rho}{c_{\Omega}}$
leads to $\kappa>12\delta^{\beta}$. On the other hand, $\rho<\left(c_{\Omega}\right)^{-\frac{\beta}{1-\beta}}$ implies that $\rho\leq\delta^{\beta}$. Hence, we obtain $[x_{1},x_{2}]\subset\Omega_{10\delta^{\beta}}$, by replacing $c_{\Omega}$ with a larger geometric constant if necessary.
In view of the above paragraph, the condition (3.1) is verified in this case. Thus, by the arguments at the beginning of this subsection, one has that
$\left|p(x_{1})-p(x_{2})\right|\leq C(d,\gamma,\Omega)\|u\|_{C^{0,\gamma}(\Omega)}^{2}|x_{1}-x_{2}|^{2\gamma}.$ (3.18)
In equations (3.4), (3.17), and (3.18) above, all the constants $C(d,\gamma,\Omega)$ depends only on $d$, $\gamma$, and the $C^{2}$-geometry of $\Omega$. Putting these estimates together, we obtain
$\frac{\left|p(x_{1})-p(x_{2})\right|}{|x_{1}-x_{2}|^{2\gamma}}\leq C(d,\gamma,\Omega)\|u\|^{2}_{C^{0,\gamma}(\Omega)}\left(1+\kappa^{-\frac{2\gamma}{\beta}}\right).$ (3.19)
This completes the proof of Theorem 1.1. We also obtain Remark 1.2 by keeping track of the dependence of constants in (3.4), (3.17), and (3.18).
@@ 4. Concluding remarks
In this note, we have given an alternative, elementary proof for the interior double Hölder regularity of the Euler pressure $p$ in bounded $C^{2}$-domains of dimension greater than or equal to $3$. We have also obtained refined characterisations for the behaviour of the quantity $|p(x_{1})-p(x_{2})|/|x_{1}-x_{2}|^{2\gamma}$ as $x_{1},x_{2}$ approaches the boundary ${\partial\Omega}$ for $\gamma\in]0,1/2[$.
We conclude with the following remarks.
@@ 4.1. Two-dimensional case
The proof of Theorem 1.1 above requires $d\geq 3$. In dimension 2, the Neumann Green function satisfies $|D^{\ell}{\mathcal{K}_{N}}(z)|\lesssim|D^{\ell}\log\,z|$ for $\ell\in\{0,1,2\}$, as opposed to the bound (3.3). An adaptation of the arguments in this note should lead to the same statement of Theorem 1.1 with $d=2$.
@@ 4.2. Unbounded domains
Acknowledgement. Both authors thank the anonymous referees for their careful reading and constructive suggestions. We are indebted to Prof. Claude Bardos, Mr. Daniel Boutros, and Prof. Edriss Titi for pointing out a flawed argument in an earlier version of the manuscript. SL also thanks Professors Linhan Li and Lihe Wang for insightful discussions on Green functions.
This research was partially supported by NSFC under Grant No. 12331008, and the Shanghai Frontier Research Institute for Modern Analysis. The research of SL was also partially supported by NSFC under Grant Nos. 12201399 and 12411530065, Young Elite Scientists Sponsorship Program by CAST 2023QNRC001, National Key Research and Development Programs 2023YFA1010900 and 2024YFA1014900, and Shanghai Rising-Star Project. The research of YGW was also partially supported by NSFC under Grant Nos. 12171317, 12250710674 and 12161141004, and Shanghai Municipal Education Commission under Grant No. 2021-01-07-00-02-E00087.
@@
Data availability statement. Our manuscript has no associated data.